# Coulomb drag driven electron-hole bifluidity in doped graphene

### v1

@@ Abstract
Motivated by the notion that a preponderance of Coulomb interactions might lead to hydrodynamics, we carry out an ab initio calculation of the charge carrier transport properties of the electron-hole plasma of doped graphene. We include both the phonon and Coulomb interactions within a momentum and band resolved Boltzmann transport formalism. We find that, under suitable conditions, the strong Coulomb drag effect induces effects like negative conductivity and joint electron-hole hydrodynamics (bifluidity) in the plasma. We also identify the exclusive electron or hole hydrodynamics. We find that there is a strong violation of the Wiedemann-Franz law in the low doped regimes. Our work elucidates the roles of the microscopic scattering mechanisms that drive these hydrodynamic phenomena.
@@ Introduction
In this work, we focus on capturing the drag effect between electrons and holes, which is the bidirectional flow of momentum between these two systems of charge carriers under non-equilibrium conditions. Since this effect is mediated by the Coulomb interaction, it also goes by the name Coulomb drag in the literature [narozhny2016coulomb]. (Generally, the Coulomb drag is defined as the drag between the charge carrier systems living on macroscopically separated materials. Graphene can be thought of as a special case where the two systems of charge carriers – electron and holes – coexist in the same material sample.) Now, in order to compute this effect within an ab initio framework, we extend the elphbolt code to include a screened Coulomb collision integral. This collision integral has been derived from first principles by Sanborn in Ref. [sanborn1995nonequilibrium] for the case of doped, 3D semiconductors, and here we simply adapt it for our 2D case. In this approximation, the exchange interaction term is ignored. In our notation, the linearized collision integral is given by
$I_{1}^{\text{Coul}}[\mathbf{J}]=$ $\dfrac{4\pi}{\hbar f^{0}_{1}(1-f^{0}_{1})A^{2}}\sum_{23}\sum_{\mathbf{G},\mathbf{G^{\prime}}}\left|\dfrac{V^{\text{2D},\infty}_{\mathbf{q}}}{\epsilon^{\text{TF}}_{\mathbf{q}}(\mathbf{G},\mathbf{G^{\prime}})}\right|^{2}$
$\delta(E_{1}+E_{2}-E_{3}-E_{4})f^{0}_{1}f^{0}_{2}(1-f^{0}_{3})(1-f^{0}_{4})$
$\left(\textbf{J}_{1}+\textbf{J}_{2}-\textbf{J}_{3}-\textbf{J}_{4}\right)\cdot\mathbf{E},$ (2)
where $\hbar$ is the reduced Planck’s constant, $E$ is the energy, $\mathbf{q}\equiv\mathbf{k}_{1}-\mathbf{k}_{3}$, $\hbar\omega\equiv E_{1}-E_{3}$, $A$ is the area of the crystal, and $V^{\text{2D},\infty}_{\mathbf{q}}$ is the 2D bare Coulomb interaction[saraga2004coulomb] screened by a high frequency dielectric. Note that the wave vector of state 4 is determined by those of the remaining three. We choose the value of $\epsilon^{\infty}$ to be that of an hBN substrate (4.14.1)[lin2013surface] to model the experimental conditions of the work in Ref. [ponomarenko2024extreme]. $\epsilon^{\text{TF}}_{\mathbf{q}}$ is the Thomas-Fermi dielectric, given by
$\epsilon_{\mathbf{q}}^{\text{TF}}(\mathbf{G},\mathbf{G}^{\prime})=\delta_{\mathbf{G},\mathbf{G}^{\prime}}+\dfrac{e^{2}\beta}{2\varepsilon_{0}A}\sum_{1}\dfrac{f^{0}_{1}(1-f^{0}_{1})}{\sqrt{|\mathbf{q}+\mathbf{G}||\mathbf{q}+\mathbf{G}^{\prime}|}},$ (3)
where $e$ is the magnitude of the electronic charge and $\varepsilon_{0}$ is the permittivity of free space. The corresponding expression for the temperature gradient field case is found through the replacement ($\mathbf{J}\rightarrow\mathbf{I}$, $\mathbf{E}\rightarrow\mathbf{\nabla}T$). Going forward, we ignore all local field effects, i.e., set $\mathbf{G}=\mathbf{G}^{\prime}$.

### v2

@@ Abstract
Motivated by the notion that a preponderance of Coulomb interactions might lead to hydrodynamics, we carry out an ab initio calculation of the charge carrier transport properties of the electron-hole gas of doped graphene. We include both the phonon and Coulomb interactions within a momentum and band resolved Boltzmann transport formalism. We find that under suitable conditions, the strong Coulomb drag effect induces phenomena like negative conductivity and joint electron-hole hydrodynamics (bifluidity) in the system. We also identify the exclusive electron or hole hydrodynamics. We find that there is a strong violation of the Wiedemann-Franz law in the low doped regimes. Our work elucidates the roles of the microscopic scattering mechanisms that drive the hydrodynamic phenomena.
@@ Introduction
In this work, we focus on capturing the drag effect between electrons and holes, which is the bidirectional flow of momentum between these two systems of charge carriers under non-equilibrium conditions. Since this effect is mediated by the Coulomb interaction, it also goes by the name Coulomb drag in the literature [narozhny2016coulomb]. (Generally, the Coulomb drag is defined as the drag between the charge carrier systems living on macroscopically separated materials. Graphene can be thought of as a special case where the two systems of charge carriers – electron and holes – coexist in the same material sample.) Now, in order to compute this effect within an ab initio framework, we extend the elphbolt code to include a screened Coulomb collision integral. This collision integral has been derived from first principles by Sanborn in Ref. [sanborn1995nonequilibrium] for the case of doped, 3D semiconductors, and here we simply adapt it for our 2D case. In this approximation, the exchange interaction term is ignored. We also ignore the local field effects. In our notation, the linearized collision integral is given by
$I_{1}^{\text{Coul}}[\mathbf{J}]=$ $\dfrac{4\pi}{\hbar f^{0}_{1}(1-f^{0}_{1})A^{2}}\sum_{23}\sum_{\mathbf{G}}|U_{3}U^{\dagger}_{1}|^{2}$
$\left|\dfrac{V^{\text{2D},\infty}_{\mathbf{q}}(\mathbf{G})}{\epsilon^{\text{TF}}_{\mathbf{q}}(\mathbf{G})}\right|^{2}\delta(E_{1}+E_{2}-E_{3}-E_{4})$
$f^{0}_{1}f^{0}_{2}(1-f^{0}_{3})(1-f^{0}_{4})\left(\textbf{J}_{1}+\textbf{J}_{2}-\textbf{J}_{3}-\textbf{J}_{4}\right)\cdot\mathbf{E},$ (2)
where $\hbar$ is the reduced Planck’s constant, $E$ is the energy, $\mathbf{q}\equiv\mathbf{k}_{1}-\mathbf{k}_{3}$, $\hbar\omega\equiv E_{1}-E_{3}$, $A$ is the area of the crystal, $V^{\text{2D},\infty}_{\mathbf{q}}$ is the 2D bare Coulomb interaction[saraga2004coulomb] screened by a high frequency dielectric, and $U^{\dagger}_{1}$ is the diagonalizer of the Wannier interpolated Hamiltonian in state $1$. Note that the wave vector of state 4 is determined by those of the remaining three. We choose the value of $\epsilon^{\infty}$ to be that of an hBN substrate (4.14.1)[lin2013surface] to model the experimental conditions of the work in Ref. [ponomarenko2024extreme]. $\epsilon^{\text{TF}}_{\mathbf{q}}$ is the Thomas-Fermi dielectric, given by
$\epsilon_{\mathbf{q}}^{\text{TF}}(\mathbf{G})=1+\dfrac{e^{2}\beta}{2\varepsilon_{0}A}\sum_{1}\dfrac{f^{0}_{1}(1-f^{0}_{1})}{|\mathbf{q}+\mathbf{G}|},$ (3)
where $e$ is the magnitude of the electronic charge and $\varepsilon_{0}$ is the permittivity of free space. The corresponding expression for the temperature gradient field case is found through the replacement ($\mathbf{J}\rightarrow\mathbf{I}$, $\mathbf{E}\rightarrow\mathbf{\nabla}T$).